# Calibrating Generative Models

### v1

@@ Abstract
Generative models frequently suffer miscalibration, wherein class probabilities and other statistics of the sampling distribution deviate from desired values. We frame calibration as a constrained optimization problem and seek the closest model in Kullback-Leibler divergence satisfying calibration constraints. To address the intractability of imposing these constraints exactly, we introduce two surrogate objectives for fine-tuning: (1) the relax loss, which replaces the constraint with a miscalibration penalty, and (2) the reward loss, which converts calibration into a reward fine-tuning problem. We demonstrate that these approaches substantially reduce calibration error across hundreds of simultaneous constraints and models with up to one billion parameters, spanning applications in protein design, image generation, and language modeling. [^1] Code: https://github.com/smithhenryd/cgm
@@ 1 Introduction
Generative models commonly produce samples whose statistics deviate systematically from desired values. Such miscalibration occurs in many domains. Image models, such as GANs and diffusion models, exhibit mode collapse, producing images that cover only a subset of the training distribution (Arora & Zhang, 2017; Qin et al., 2023). Language models represent gender, race, religion, and age in ways that reinforce societal biases (Gallegos et al., 2024). In synthetic biology applications, protein structure models produce samples that have alpha-helical and beta-strand substructures at frequencies atypical of proteins found in nature (Lu et al., 2025), and DNA models generate samples that contain subsequences at frequencies that differ from those in human DNA (Sarkar et al., 2024). These calibration errors arise from many sources including dataset imbalances, suboptimal training dynamics, and post-hoc adjustments such as low-temperature sampling or preference fine-tuning.
@@
Related work. Within the generative modeling community, there has been a wealth of fine-tuning methods that incorporate preferences at the level of individual samples through a user-specified reward (Christiano et al., 2017; Rafailov et al., 2023; Uehara et al., 2024; Domingo-Enrich et al., 2025). None of these methods solves the calibration problem, which imposes a hard constraint at the distribution level.
Two prior works (Khalifa et al., 2021; Shen et al., 2024) have proposed fine-tuning procedures for distribution level constraints, but each applies to a single class of generative models. Khalifa et al. (2021), the most similar to the present work, fine-tunes autoregressive language models to match distributional constraints with an algorithm similar to CGM-reward. Shen et al. (2024) propose a method for balancing class proportions in diffusion models that relies upon optimal transport. Compared to the present work, neither method reduces a majority of calibration error, and Khalifa et al. (2021) demonstrates their algorithm only for low-dimensional (<10{<}10) constraints. By contrast we show CGM reduces the majority of calibration error in all applications and is successful at calibrating to $10^{2}$ constraints. Appendix˜A provides an extended discussion of related work.
@@ 2.1 The Relax Loss
where $\lambda>0$ is a hyperparameter that trades off between satisfying the calibration constraint and minimizing the KL divergence. In our experiments we choose $\lambda$ by a grid-search. In the limit as $\lambda\to 0$, $\mathcal{L}^{\text{viol}}$ is the dominant term in the relax loss, and we expect the minimizer of (2) to approach the solution of the calibration problem (1).
Suppose we have $M$ independent samples $\{\boldsymbol{x}_{m}\}_{m=1}^{M}$ from $p_{\theta}$. To estimate $\mathcal{L}^{\text{relax}}$ we separately estimate the KL divergence ($\mathcal{L}^{\text{KL}}$) by
$\widehat{\mathcal{L}}^{\text{KL}}:=\frac{1}{M}\sum_{m=1}^{M}\log\frac{p_{\theta}(\boldsymbol{x}_{m})}{p_{\theta_{\text{base}}}(\boldsymbol{x}_{m})},$
and the constraint penalty ($\mathcal{L}^{\text{viol}}$) by
@@
where the second term is a bias correction. Combining these estimators yields our overall estimator for the relax objective, $\widehat{\mathcal{L}}^{\text{relax}}=\widehat{\mathcal{L}}^{\text{viol}}+\lambda\widehat{\mathcal{L}}^{\text{KL}}$. Section˜B.1 shows $\widehat{\mathcal{L}}^{\text{relax}}$ is unbiased for $\mathcal{L}^{\text{relax}}$.
@@ 2.2 The Reward Loss
The reward loss avoids the intractability of imposing the calibration constraint exactly by leveraging a connection between the calibration problem (1) and the maximum entropy problem (Jaynes, 1957; Kullback, 1959; Csiszár, 1975). We first introduce the maximum entropy problem and then show how to approximate its solution with samples from $p_{\theta_{\text{base}}}$. Lastly, we propose the reward loss as a divergence to this approximate solution and describe connections to reward fine-tuning.
@@
$\underset{p\in\mathcal{P}(p_{\theta_{\text{base}}})}{\operatorname*{arg\,min}}\textrm{D}_{\textrm{KL}}\left(p\;\|\;p_{\theta_{\text{base}}}\right),\quad\text{such that}~\mathbb{E}_{p}[\boldsymbol{h}(\boldsymbol{x})]=\boldsymbol{h}^{\ast}$ (4)
@@
Under assumptions, there exists a unique solution to (4) that has the form
$p_{\boldsymbol{\alpha}^{\ast}}(\boldsymbol{x})\propto p_{\theta_{\text{base}}}(\boldsymbol{x})\exp\left\{r_{\boldsymbol{\alpha}^{\ast}}(\boldsymbol{x})\right\},\quad r_{\boldsymbol{\alpha}}(\boldsymbol{x}):=\boldsymbol{\alpha}^{\top}\boldsymbol{h}(\boldsymbol{x}).$ (5)
Appendix˜C provides further exposition to the maximum entropy problem as well as a proof.
@@
In Section˜C.3 we demonstrate that a similar statement holds for the relax loss: when $p_{\theta}$ is sufficiently expressive, the optimum of the relax loss is close to $p_{\lambda}(\boldsymbol{x})\propto p_{\theta_{\text{base}}}(\boldsymbol{x})\exp\left\{r_{\boldsymbol{\alpha}_{\lambda}}(\boldsymbol{x})\right\}$, where $\boldsymbol{\alpha}_{\lambda}$ depends on the regularization strength $\lambda>0$ and is not generally equal to $\boldsymbol{\alpha}^{\ast}$. However, as $\lambda\to 0$, $\|\boldsymbol{\alpha}_{\lambda}-\boldsymbol{\alpha}^{\ast}\|$ approaches zero at rate $\lambda$. This formalizes our intuition from Section˜2.1 that as $\lambda\to 0$, the relax loss solves the calibration problem.
Estimating $p_{\boldsymbol{\alpha}^{\ast}}$. The idea of minimizing a divergence to $p_{\boldsymbol{\alpha}^{\ast}}$ introduces a challenge: even when the solution $p_{\boldsymbol{\alpha}^{\ast}}$ to the maximum entropy problem (4) exists, its parameters $\boldsymbol{\alpha}^{\ast}$ are not immediately computable. To address this challenge, we leverage Wainwright & Jordan (2008, Theorem 3.4), which states that, under conditions, solving problem (4) is equivalent to computing
$\operatorname*{arg\,max}_{\boldsymbol{\alpha}}\boldsymbol{\alpha}^{\top}\boldsymbol{h}^{\ast}-\log\left(\int\exp\{r_{\boldsymbol{\alpha}}(\boldsymbol{x})\}p_{\theta_{\text{base}}}(\boldsymbol{x})d\boldsymbol{x}\right).$ (6)
@@
$p_{\theta_{\mathrm{base}}},h(\cdot),\boldsymbol{h}^{\ast},M,$ and $\lambda$
@@
$p_{\theta_{\mathrm{base}}},h(\cdot),\boldsymbol{h}^{\ast},M,N$
@@ 2.3 Gradient Estimation
We next describe our approach to computing unbiased estimates for the gradients of $\mathcal{L}^{\text{relax}}(\theta)$ and $\mathcal{L}^{\text{reward}}(\theta)$. This enables optimization of the relax and reward losses via stochastic optimization. We leverage the score function gradient estimator (Williams, 1992; Ranganath et al., 2014) and a similar importance sampling-based gradient estimator for the relax loss.
@@
for any set of model parameters $\theta^{\prime}$. Since the expectation in equation˜9 does not depend on $\theta,$ we can approximate its gradient with Monte Carlo samples from $p_{\theta^{\prime}}$. Moreover, since the choice of $\theta^{\prime}$ is arbitrary, we can choose $\theta^{\prime}{=}\theta$, draw independent samples $\boldsymbol{x}_{m}$ from $p_{\theta}$, and compute
$\left[\nabla_{\theta}\mathcal{L}(\theta,\theta^{\prime})\right]~|~_{\theta^{\prime}=\theta}\approx\frac{1}{M}\sum\frac{1}{p_{\theta}(\boldsymbol{x}_{m})}\nabla_{\theta}[p_{\theta}(\boldsymbol{x}_{m})f(\boldsymbol{x}_{m},\theta)].$ (10)
@@
Although score function gradient estimates are known to suffer from high variance (Mohamed et al., 2020), we show that, paired with variance reduction strategies (Section˜B.2), they perform well even in problem settings with high-dimensional latent variables, such as diffusion models and masked language models (Section˜4.1).
Figure 1: Calibrating mixture proportions in a 1D GMM. A: The CGM-relax and CGM-reward solutions closely approximate the maximum entropy solution. B: (top) The CGM-relax regularization parameter $\lambda$ trades off between constraint satisfaction and closeness to the base model (bottom) CGM-reward is accurate when enough samples $N$ are used to estimate $\boldsymbol{\alpha}^{\ast}$.
@@ 3 Simulations: Determining when CGM thrives and struggles
To understand the success and failure cases of CGM, we perform evaluations in a tractable “toy” setting. This setting allows us to understand the role of the CGM hyperparameters $\lambda$ and $N$, and to test CGM in challenging problem settings, including rare events and high-dimensional constraints.
Simulation setup and evaluation. We consider fine-tuning a diffusion model targeting a Gaussian mixture model (GMM) to reweight the mixture proportions of each mode. Here, $p_{\theta}(\boldsymbol{x})$ is a generative model of continuous paths $\boldsymbol{x}=(\boldsymbol{x}(t))_{t\in[0,1]}$, whose evolution is described by a stochastic differential equation (SDE). To sample from $p_{\theta}$, one first draws $\boldsymbol{x}(0)$ from the tractable initial distribution and then simulates the SDE starting from time t=0t{=}0 up until time t=1t{=}1. Section˜D.1 provides additional background on diffusion models.
@@
Selecting hyperparameters for CGM-relax and CGM-reward. We first initialize our base model $p_{\theta_{\text{base}}}$ such that $p_{\theta_{\text{base}}}(\boldsymbol{x}(1))$ is a one-dimensional Gaussian mixture with two well separated modes (Figure˜1 A). We define the calibration problem with statistic $\boldsymbol{h}(\boldsymbol{x})=\mathbbm{1}\{\boldsymbol{x}(1)>0\}$ to upweight the mass in right mode from $\mathbb{E}_{p_{\theta_{\text{base}}}}[\boldsymbol{h}(\boldsymbol{x}(1))]=0.5$ to $\boldsymbol{h}^{\ast}=0.8$.
@@
For CGM-relax we observe that the regularization parameter $\lambda$ trades off between constraint satisfaction and deviation from the base model (Figure˜1B). With large $\lambda$ the model deviates little from $p_{\theta_{\text{base}}}$ but does not satisfy the constraint, whereas for small $\lambda$ the model satisfies the constraint but has KL to $p_{\theta_{\text{base}}}$ that exceeds that of the maximum entropy solution. For CGM-reward, we observe that increasing $N$ results in more accurate recovery of the variational parameters $\boldsymbol{\alpha}^{\ast}$ and thereby a better approximation to the maximum entropy solution. For appropriate hyperparameters, both solve the calibration problem to high accuracy. In the remaining experiments, we perform grid-search to select $\lambda$ in CGM-relax and use $N=10^{5}$ samples to estimate $\boldsymbol{\alpha}^{\ast}$ in CGM-reward.
@@
Upweighting rare events. Increasing the proportion of generations belonging to rare classes is central to applications including protein ensemble modeling (Lewis et al., 2025) and reinforcement learning (O’Kelly et al., 2018). To assess the performance of CGM in this setting, we consider variations of the GMM reweighting problem in which we consider increasingly small mixture proportions $\pi=\mathbb{E}_{p_{\theta_{\text{base}}}}[\boldsymbol{h}(\boldsymbol{x}(1))]$ of the mode to upweight by calibration. We vary $\pi$ from $\boldsymbol{h}^{\ast}=0.8$ (already calibrated) to $10^{-5}$ and use a constant batch size $M=10^{2}$.
@@
We find that both algorithms perform well with base model event rarity as small as $\pi=10^{-3}$; the majority of miscalibration is reduced without divergence from the base model much larger than the maximum entropy solution (Figure˜2A). This is surprising since for $\pi=10^{-3}$, most batches sampled from $p_{\theta_{\text{base}}}$ contain no samples belonging to the second mode. Performance degrades below this threshold, but we suspect larger batch sizes would allow upweighting even rarer events.
@@
Scalability to high-dimensional models and constraints. We next evaluate how performance depends on the dimensionality, $k$, of the GMM and the constraint. We take the base model to be a product of one-dimensional GMMs with marginals as in Figure˜1A. For the calibration constraint, we choose the $\boldsymbol{h}(\boldsymbol{x})=[\mathbbm{1}\{\boldsymbol{x}(1)[1]>0\},\dots,\mathbbm{1}\{\boldsymbol{x}(1)[k]>0\}]$, where $\boldsymbol{x}(1)[i]$ is the $i$th dimension of $\boldsymbol{x}(1)$ and $\boldsymbol{h}^{\ast}=[0.8,\dots,0.8]$. Since both the base model $p_{\theta_{\text{base}}}$ and maximum entropy solution $p_{\boldsymbol{\alpha}^{\ast}}$ are independent across dimension, the KL distance between these two distributions grows linearly in dimension. The multimodality of this model, with $2^{k}$ modes, mimics the multimodality of practical generative models. We perform CGM-relax and CGM-reward with batch size $M=10^{4}$.
@@
In this high-dimensional regime, significant discrepancies emerge between CGM-relax and CGM-reward (Figure˜2B). CGM-relax consistently eliminates the majority of constraint violation up to $k{=}10^{3}$, albeit with a non-trivial excess KL divergence to $p_{\theta_{\text{base}}}$ compared to the maximum entropy solution $p_{\boldsymbol{\alpha}^{*}}$ that increases linearly with dimension. Although CGM-reward performs well for low-dimensional constraints (<10{<}10), we find that the empirical maximum entropy problem (7) is infeasible with high probability for >30{>}30 constraints. In fact, even when $\widehat{\boldsymbol{\alpha}}_{N}$ is fixed to its oracle value $\boldsymbol{\alpha}^{\ast}$ (Figure˜2B), CGM-relax still outperforms CGM-reward.

### v2

@@ Abstract
Generative models frequently suffer miscalibration, wherein statistics of the sampling distribution such as class probabilities deviate from desired values. We frame calibration as a constrained optimization problem and seek the closest model in Kullback-Leibler divergence satisfying calibration constraints. To address the intractability of imposing these constraints exactly, we introduce two surrogate objectives for fine-tuning: (1) the relax loss, which replaces the constraint with a miscalibration penalty, and (2) the reward loss, which converts calibration into a reward fine-tuning problem. We demonstrate that these approaches substantially reduce calibration error across hundreds of simultaneous constraints and models with up to nine billion parameters, spanning applications in protein design, image generation, and language modeling. [^1] Code: https://github.com/smithhenryd/cgm
@@ 1 Introduction
Generative models commonly produce samples whose statistics deviate systematically from desired values. Such miscalibration occurs in many domains. Image models, such as GANs and diffusion models, exhibit mode collapse, producing images that cover only a subset of the training distribution (Arora and Zhang, 2017; Qin et al., 2023). Language models represent gender, race, religion, and age in ways that reinforce societal biases (Gallegos et al., 2024). In synthetic biology applications, protein structure models produce samples that have alpha-helical and beta-strand substructures at frequencies atypical of proteins found in nature (Lu et al., 2025), and DNA models generate samples that contain subsequences at frequencies that differ from those in human DNA (Sarkar et al., 2024). These calibration errors arise from many sources including dataset imbalances, suboptimal training dynamics, and post-hoc adjustments such as low-temperature sampling or preference fine-tuning.
@@
Related work. Beyond the context of generative modeling, calibration is a major topic in supervised machine-learning (e.g., Lichtenstein et al., 1977; Dawid, 1982; Naeini et al., 2015; Guo et al., 2017; Vaicenavicius et al., 2019). In these settings, a model is considered calibrated when events predicted to occur with a particular probability occur at that rate on future data. This notion of calibration may also be expressed an expectation constraint, and variety of methods exist to adapt the outputs of a predictive models to achieve calibration such as Platt scaling (Platt, 1999) or conformal methods (Shafer and Vovk, 2008) for more general prediction sets. However, these approaches do not apply our setting, where the goal is to alter the sampling distribution of the generative model.
Figure 1: CGM aims to identify the generative model $p_{\theta^{*}}$, among those that satisfy the moment constraint, that is closest to $p_{\theta_{\text{base}}}$. $p_{\theta^{*}}$ is related to the maximum entropy distribution $p_{\boldsymbol{\alpha}^{*}}$.
Within the generative modeling community, there are a wealth of fine-tuning methods that incorporate preferences at the level of individual samples through a user-specified reward (Christiano et al., 2017; Rafailov et al., 2023; Uehara et al., 2024; Domingo-Enrich et al., 2025). These methods do not address calibration, for which the goal is to impose a constraint on the distribution $p_{\theta}$ rather than its samples $\boldsymbol{x}$.
Two prior works (Khalifa et al., 2021; Shen et al., 2024) propose fine-tuning procedures for distribution level constraints, but each applies to a single model class. Khalifa et al. (2021), the most similar to the present work, fine-tunes autoregressive language models to match distributional constraints with an algorithm similar to CGM-reward. Shen et al. (2024) propose a method for balancing class proportions in diffusion models that relies upon optimal transport. Compared to the present work, neither method reduces a majority of calibration error.
@@ 2.1 The Relax Loss
where $\lambda>0$ is a hyperparameter. In the limit as $\lambda\to 0$, $\mathcal{L}^{\text{viol}}$ is the dominant term in the relax loss, and we expect the minimizer of (2) to approach the solution of the calibration problem (1). In practice, finite $\lambda$ must be chosen, and in our experiments we find $\lambda$ trades off between satisfying the calibration constraint and remaining close to $p_{\theta_{\text{base}}}$. Section˜3 suggests a heuristic for this choice.
We choose the constraint penalty $\mathcal{L}^{\text{viol}}$ to be a squared $\ell_{2}$ norm because it is amenable to unbiased estimation. Suppose we have $M$ independent samples $\{\boldsymbol{x}_{m}\}_{m=1}^{M}$ from $p_{\theta}$. An unbiased estimate of $\mathcal{L}^{\text{viol}}$ is then
@@
where the second term is a bias correction.
Similarly, an unbiased estimate of the KL divergence ($\mathcal{L}^{\text{KL}}$) is
$\widehat{\mathcal{L}}^{\text{KL}}:=\frac{1}{M}\sum_{m=1}^{M}\log\frac{p_{\theta}(\boldsymbol{x}_{m})}{p_{\theta_{\text{base}}}(\boldsymbol{x}_{m})}.$
Combining these estimators yields our overall estimator for the relax objective, $\widehat{\mathcal{L}}^{\text{relax}}=\widehat{\mathcal{L}}^{\text{viol}}+\lambda\widehat{\mathcal{L}}^{\text{KL}}$. Section˜B.1 shows $\widehat{\mathcal{L}}^{\text{relax}}$ is unbiased for $\mathcal{L}^{\text{relax}}$.
@@ 2.2 The Reward Loss
The reward loss avoids the intractability of imposing the calibration constraint exactly through a connection between the calibration problem and the maximum entropy problem (Jaynes, 1957; Kullback, 1959; Csiszár, 1975). We first introduce the maximum entropy problem and then show how to approximate its solution with samples from $p_{\theta_{\text{base}}}$. Lastly, we propose the reward loss as a divergence to this approximate solution and describe connections to reward fine-tuning.
@@
$\underset{p\in\mathcal{P}(p_{\theta_{\text{base}}})}{\operatorname*{arg\,min}}\textrm{D}_{\textrm{KL}}\left(p\;\|\;p_{\theta_{\text{base}}}\right),\quad\text{such that}~\mathbb{E}_{p}[\boldsymbol{h}(\boldsymbol{x})]=\boldsymbol{h}^{\ast},$ (4)
@@
Suppose $\boldsymbol{h}^{\ast}$ lies in the relative interior of the set of moments of $\boldsymbol{h}$ attainable by distributions $P\in\mathcal{P}(p_{\theta_{\text{base}}})$. Then there exists a vector $\boldsymbol{\alpha}^{\ast}$ for which
$p_{\boldsymbol{\alpha}^{\ast}}(\boldsymbol{x})\propto p_{\theta_{\text{base}}}(\boldsymbol{x})\exp\left\{r_{\boldsymbol{\alpha}^{\ast}}(\boldsymbol{x})\right\},\quad r_{\boldsymbol{\alpha}}(\boldsymbol{x}):=\boldsymbol{\alpha}^{\top}\boldsymbol{h}(\boldsymbol{x})$ (5)
satisfies $\mathbb{E}_{p_{\boldsymbol{\alpha}^{\ast}}}[\boldsymbol{h}(\boldsymbol{x})]=\boldsymbol{h}^{\ast}$ and $p_{\boldsymbol{\alpha}^{\ast}}$ is the solution to the maximum entropy problem.
Appendix˜C provides additional background on the maximum entropy problem and a proof.
@@
Estimating $p_{\boldsymbol{\alpha}^{\ast}}$. The idea of minimizing a divergence to $p_{\boldsymbol{\alpha}^{\ast}}$ introduces a challenge: although Theorem˜2.1 establishes the existence of some $\boldsymbol{\alpha}^{\ast}$ for which $p_{\boldsymbol{\alpha}^{\ast}}$ solves the maximum entropy problem, it does not tell us how to compute it. To address this challenge, we leverage Wainwright and Jordan (2008, Theorem 3.4), which states that when the assumption of Theorem˜2.1 holds and there are no redundancies among the constraints $\boldsymbol{h}$
$\boldsymbol{\alpha}^{\ast}=\operatorname*{arg\,max}_{\boldsymbol{\alpha}}\boldsymbol{\alpha}^{\top}\boldsymbol{h}^{\ast}-\log\left(\int\exp\{r_{\boldsymbol{\alpha}}(\boldsymbol{x})\}p_{\theta_{\text{base}}}(\boldsymbol{x})d\boldsymbol{x}\right).$ (6)
@@
$p_{\theta_{\mathrm{base}}},\boldsymbol{h}(\cdot),\boldsymbol{h}^{\ast},M,$ and $\lambda$
@@
$p_{\theta_{\mathrm{base}}},\boldsymbol{h}(\cdot),\boldsymbol{h}^{\ast},M,N$
@@ 2.3 Gradient Estimation
We next describe our approach to computing unbiased estimates for the gradients of $\mathcal{L}^{\text{relax}}(\theta)$ and $\mathcal{L}^{\text{reward}}(\theta)$. This enables optimization of the relax and reward losses by stochastic optimization. We leverage the score function gradient estimator (Williams, 1992; Ranganath et al., 2014) and a similar importance sampling-based gradient estimator for the relax loss.
@@
for any set of model parameters $\theta^{\prime}$. Since the expectation in equation˜9 does not depend on $\theta,$ we can approximate its gradient with Monte Carlo samples from $p_{\theta^{\prime}}$. The density ratio $p_{\theta}(\boldsymbol{x}_{m})/p_{\theta^{\prime}}(\boldsymbol{x}_{m})$ in equation˜9 can be understood as the weights of an importance sampling estimate against target $p_{\theta}$ with proposal $p_{\theta^{\prime}}$.
@@
Although score function gradient estimates are known to suffer from high variance (Mohamed et al., 2020), we show that, paired with the variance reduction strategies described in Section˜B.2, they perform well even in problem settings with high-dimensional latent variables, such as diffusion models and masked language models (Section˜4.1).
Figure 2: Calibrating mixture proportions in a diffusion model targeting a 1D GMM. A: The CGM-relax and CGM-reward solutions closely approximate the maximum entropy solution. B: (top) The CGM-relax regularization parameter $\lambda$ trades off between constraint satisfaction and closeness to the base model (bottom) CGM-reward is accurate when enough samples $N$ are used to estimate $\boldsymbol{\alpha}^{\ast}$.
2.4 Relationship Between Relax and Reward Losses
We showed in Section˜2.2 that when the generative model $p_{\theta}$ is replaced by a probability distribution $p\in\mathcal{P}(p_{\theta_{\text{base}}})$, the minimizer of the reward loss is equal to the base model $p_{\theta_{\text{base}}}$ tilted by the exponentiated reward with parameters $\boldsymbol{\alpha}=\boldsymbol{\alpha}^{\ast}$.
We demonstrate in Section˜C.3 that a similar statement holds for the relax loss: when $p_{\theta}$ is replaced by $p\in\mathcal{P}(p_{\theta_{\text{base}}})$, the minimizer of the relax loss is equal to $p_{\lambda}(\boldsymbol{x})\propto p_{\theta_{\text{base}}}(\boldsymbol{x})\exp\left\{r_{\boldsymbol{\alpha}_{\lambda}}(\boldsymbol{x})\right\}$, where the vector $\boldsymbol{\alpha}_{\lambda}$ depends on the regularization strength $\lambda>0$ and is not generally equal to $\boldsymbol{\alpha}^{\ast}$. However, we show in Section˜C.3 that as $\lambda\to 0$, both $\|\boldsymbol{\alpha}_{\lambda}-\boldsymbol{\alpha}^{\ast}\|$ and the miscalibration error $\|\mathbb{E}_{q_{\lambda}}[\boldsymbol{h}(\boldsymbol{x})]-\boldsymbol{h}^{\ast}\|$ approach zero at rate $\mathcal{O}(\lambda)$ (Hestenes, 1969). Moreover, up to first order, the miscalibration of $p_{\lambda}$ depends on the curvature of $p_{\boldsymbol{\alpha}}(\boldsymbol{x})\propto p_{\theta_{\text{base}}}(\boldsymbol{x})\exp\{r_{\boldsymbol{\alpha}}(\boldsymbol{x})\}$ about $\boldsymbol{\alpha}=\boldsymbol{\alpha}^{\ast}$, as measured by the minimum eigenvalue of the Fisher information matrix.
@@ 3 Simulations: Determining when CGM thrives and struggles
To understand the success and failure cases of CGM, we perform evaluations in a tractable setting. This setting allows us to understand the role of the CGM hyperparameters $\lambda$ and $N$, and to test CGM in challenging cases, including rare events and high-dimensional constraints. We provide an overview of continuous-time diffusion models in Appendix˜D and further discussion of experimental details in Section˜E.1.
Simulation setup and evaluation. We consider fine-tuning a diffusion model targeting a Gaussian mixture model (GMM) to reweight the mixture proportions of each mode. Here, $p_{\theta}(\boldsymbol{x})$ is a generative model of continuous paths $\boldsymbol{x}=(\boldsymbol{x}(t))_{t\in[0,1]}$, whose evolution is described by a stochastic differential equation (SDE). To sample from $p_{\theta}$, one first draws $\boldsymbol{x}(0)$ from the tractable initial distribution and then simulates the SDE starting from time t=0t{=}0 up until time t=1t{=}1.
@@
Selecting hyperparameters for CGM-relax and CGM-reward. We first initialize our base model $p_{\theta_{\text{base}}}$ such that $p_{\theta_{\text{base}}}(\boldsymbol{x}(1))$ is a one-dimensional Gaussian mixture with two well separated modes (Figure˜2A). We define the calibration problem with statistic $\boldsymbol{h}(\boldsymbol{x})=\mathbbm{1}\{\boldsymbol{x}(1)>0\}$ to upweight the mass in right mode from $\mathbb{E}_{p_{\theta_{\text{base}}}}[\boldsymbol{h}(\boldsymbol{x}(1))]=0.5$ to $\boldsymbol{h}^{\ast}=0.8$.
@@
For CGM-relax we observe that the regularization parameter $\lambda$ trades off between constraint satisfaction and deviation from the base model (Figure˜2B). With large $\lambda$ the model deviates little from $p_{\theta_{\text{base}}}$ but does not satisfy the constraint, whereas for small $\lambda$ the model satisfies the constraint but has KL to $p_{\theta_{\text{base}}}$ that exceeds that of the maximum entropy solution. For CGM-reward, we observe that increasing $N$ results in more accurate recovery of the variational parameters $\boldsymbol{\alpha}^{\ast}$ and thereby a better approximation to the maximum entropy solution. For appropriate hyperparameters, both solve the calibration problem to high accuracy.
In the remaining experiments, we adopt the following heuristic for selecting $\lambda$ in CGM-relax: we choose $\log(\lambda)$ on a $10$-point linear grid defined over the interval [3,0][-3,0] and perform calibration for each value. Among these values, we select the largest $\lambda$ for which constraint violation is reduced relative to the base model by 90%. For CGM-reward, we use $N=10^{5}$ samples to estimate $\boldsymbol{\alpha}^{\ast}$.
@@
Upweighting rare events. Increasing the proportion of generations belonging to rare classes is central to applications including protein ensemble modeling (Lewis et al., 2025) and reinforcement learning (O’Kelly et al., 2018). To assess the performance of CGM in this setting, we consider variations of the GMM reweighting problem in which we consider increasingly small $p_{\theta_{\text{base}}}(\boldsymbol{x}(1)>0)=\mathbb{E}_{p_{\theta_{\text{base}}}}[\boldsymbol{h}(\boldsymbol{x}(1))]$. When $p_{\theta_{\text{base}}}(\boldsymbol{x}(1)>0)$ is greater than $10^{-3}$, it is approximately equal to the mixture proportion of the mode to upweight by calibration. We vary $p_{\theta_{\text{base}}}(\boldsymbol{x}(1)>0)$ from $\boldsymbol{h}^{\ast}=0.8$ (already calibrated) to approximately $10^{-5}$ and use a constant batch size $M=10^{2}$.
@@
We find that both algorithms perform well with base model event rarity as small as $10^{-3}$; the majority of miscalibration is reduced without divergence from the base model much larger than the maximum entropy solution (Figure˜3A). This is surprising since for $10^{-3}$, most batches sampled from $p_{\theta_{\text{base}}}$ contain no samples belonging to the second mode. Performance degrades below this threshold, but we suspect larger batch sizes would allow upweighting even rarer events.
@@
Scalability to high-dimensional models and constraints. We next evaluate how performance depends on the dimensionality, $k$, of the GMM and the constraint. We take the base model to be a product of one-dimensional GMMs with marginals as in Figure˜2A. For the calibration constraint, we choose the $\boldsymbol{h}(\boldsymbol{x})=[\mathbbm{1}\{\boldsymbol{x}(1)[1]>0\},\dots,\mathbbm{1}\{\boldsymbol{x}(1)[k]>0\}]$, where $\boldsymbol{x}(1)[i]$ is the $i$th dimension of $\boldsymbol{x}(1)$ and $\boldsymbol{h}^{\ast}=[0.8,\dots,0.8]$. Since both the base model $p_{\theta_{\text{base}}}$ and maximum entropy solution $p_{\boldsymbol{\alpha}^{\ast}}$ are independent across dimension, the KL distance between these two distributions grows linearly in dimension. The multimodality of this model, with $2^{k}$ modes, mimics the multimodality of practical generative models. We perform CGM-relax and CGM-reward with batch size $M=10^{4}$.
@@
In this high-dimensional regime, significant discrepancies emerge between CGM-relax and CGM-reward (Figure˜3B). CGM-relax consistently eliminates the majority of constraint violation up to $k{=}10^{3}$, albeit with a non-trivial excess KL divergence to $p_{\theta_{\text{base}}}$ compared to the maximum entropy solution $p_{\boldsymbol{\alpha}^{*}}$ that increases linearly with dimension. Although CGM-reward performs well for low-dimensional constraints (<10{<}10), we find that the empirical maximum entropy problem (7) is infeasible with high probability for >30{>}30 constraints. In fact, even when $\widehat{\boldsymbol{\alpha}}_{N}$ is fixed to its oracle value $\boldsymbol{\alpha}^{\ast}$ (Figure˜3B), CGM-relax still outperforms CGM-reward.